OpenStudy (anonymous):

if d/dx = (f(2x^5))=6x^5 calculate f'(x)

OpenStudy (loser66):

\[\frac{d}{dx}=(f(2x^5))=6x^5\]I don't understand, please explain

OpenStudy (mertsj):

I don't know. Have to ask the smart people. @Phi

OpenStudy (phi):

any other clues ?

OpenStudy (loser66):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

d/dx[ f(2x^5) ] = f ' (2x^5) * d/dx[2x^5] d/dx[ f(2x^5) ] = f ' (2x^5) * 10x^4 -------------------------------------- d/dx[ f(2x^5) ] = 6x^5 f ' (2x^5) * 10x^4 = 6x^5 f ' (2x^5) = 6x^5/10x^4 f ' (2x^5) = (3/5)x that's as far as I could get

OpenStudy (anonymous):

so the answer is (3/5)x?

jimthompson5910 (jim_thompson5910):

well no because they want f ' (x)

jimthompson5910 (jim_thompson5910):

but not sure what the function f ' (x) could be, something seems missing, but idk

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np, i would ask your teacher about it

OpenStudy (phi):

ok, that makes sense @jim_thompson5910 how about a change of variables? f ' (2x^5) = (3/5)x let u= 2x^5 and x= (u/2)^(1/5) f'(u) = (3/5) (u/2)^(1/5) or , renaming u to x \[ f'(x)= \frac{3}{5} \sqrt[5]{\frac{x}{2}} \]

OpenStudy (loser66):

I am with phi, then